# First-Order Sparse Convex Optimization: Better Rates with Sparse Updates

### v2

@@ 2 Unified Setup, Algorithm, and Convergence Analysis
Consider the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ of Algorithm 1 with step-size $\eta=\frac{\alpha_{2}}{4\beta(4s+s^{*}_{\mathcal{K}})}$. Then,
$\forall t\geq 1:\quad f({\mathbf{x}}_{t+1})-f^{*}\leq\left({f({\mathbf{x}}_{t})-f^{*}}\right)\left({1-\frac{\alpha_{2}}{8\beta(4s+s^{*}_{\mathcal{K}})}}\right).$
Proof.
Fix some iteration $t$ of the algorithm. The Lipchitz continuity of the gradient implies the following well known inequality:
@@
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
@@
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Using the definition of ${\mathbf{v}}_{t}$ in the algorithm and the assumption that ${\mathbf{x}}^{*}\in\mathcal{K}_{s}$ we have that,
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Using the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm, Item 3 and then Item 2 in Assumption 1 we have that,
$\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}\leq\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|^{2}\leq s_{\mathcal{K}}^{*}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}.$
@@
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}s^{*}_{\mathcal{K}}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(2s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}+2\beta\eta^{2}s\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|_{2}^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2},$
where the last inequality follows from the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm.
@@
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})-\eta\left({f({\mathbf{x}}_{t})-f({\mathbf{x}}^{*})}\right)+\frac{2\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})}{\alpha_{2}}\left({f({\mathbf{x}}_{t})-f({\mathbf{x}}^{*})}\right).$
Subtracting $f({\mathbf{x}}^{*})$ from both sides and plugging-in our choice of step-size $\eta$, completes the proof.
@@ 3.1 $\ell_{1}$-ball setup
[…]
We shall require the following simple lemma (a proof is brought in the appendix for completeness).
@@
$\|{{\mathbf{x}}-{\mathbf{y}}}\|_{1}\leq 2\sqrt{\textrm{nnz}({\mathbf{x}})}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
Let $s$ be such that $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\textrm{nnz}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\|{{\mathbf{x}}}\|_{1}\leq R,~\textrm{nnz}({\mathbf{x}})\leq s\}$. In light of Lemma 1 and the assumption that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=2s$.
@@
Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, by computing the Euclidean projection of the point ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$ onto the $s$-sparsified $\ell_{1}$-ball $\mathcal{K}_{s}$, is done by computing the Euclidean projection of the vector in $\mathbb{R}^{s}$ which corresponds to the $s$-largest (in absolute value) entries in ${\mathbf{z}}_{t}$ onto the $s$-dimensional $\ell_{1}$-ball of radius $R$, and then appending zeros (in the entries which correspond to the $n-s$ lower entries in ${\mathbf{z}}_{t}$) to obtain again a point the $n$-dimensional $\ell_{1}$-ball of radius $R$. See [3] for proof of the correctness of this procedure.
@@
When applied to Problem (7) with an optimal solution ${\mathbf{x}}^{*}$ which satisfies $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside of gradient computations, the runtime of each iteration is $O(n\log{s})$. In particular, if $f$ is (convex) quadratic, the overall runtime of each iteration is $O(sn)$.
@@
It is straightforward to extend the result above to the case in which ${\mathbf{x}}$ is constrained to be in $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{x}}~|~\|{{\mathbf{x}}}\|_{1}\leq R\}$, where $\mathcal{C}\subset\mathbb{R}^{n}$ is closed and convex, as long as ${\mathbf{x}}^{*}$ satisfies $\|{{\mathbf{x}}}\|_{1}=R$. For instance when $\mathcal{C}=\{{\mathbf{x}}~|~{\mathbf{x}}\geq 0,\sum_{i=1}^{n}{\mathbf{x}}_{i}=R\}$, we have that $\mathcal{K}$ is simply a scaled simplex for which the assumption $\|{{\mathbf{x}}^{*}}\|_{1}=R$ always holds. Moreover, when $\mathcal{C}$ is closed under permutation of coordinates (e.g., an $\ell_{p}$ ball or its non-negative restriction), the sparse proximal step (computation of ${\mathbf{v}}_{t}$ in Algorithm 1) admits efficient implementation as detailed in [3].
3.2 Bounded positive semidefinite cone setup
@@
$\min_{{\mathbf{X}}\in\mathbb{S}^{n}:~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R}f({\mathbf{X}}).$ (8)
[…]
The following lemma is analogues to Lemma 1. A proof is given in the appendix for completeness.
@@
Let ${\mathbf{X}},{\mathbf{Y}}\in\{{\mathbf{Z}}\in\mathbb{S}^{n}~|~{\mathbf{Z}}\succeq 0\}$ such that ${\mathbf{X}}\neq\mathbf{0}$, $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$. Then,
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq 4\sqrt{\textrm{rank}({\mathbf{X}})}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}.$
Let $s$ be such that $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~\textrm{rank}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\|{{\mathbf{X}}}\|_{1}\leq R,~\textrm{rank}({\mathbf{X}})\leq s\}$. In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=16s$.
[…]
Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, is done as follows: compute the $s$ top components (corresponding to the largest signed eigenvalues) in the eigen-decomposition of the matrix ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$, and denote it as $\sum_{i=1}^{s}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, ${\mathbf{u}}_{i}\in\mathbb{R}^{n},i=1,\dots,s$. Then, project the vector of eigenvalues $(\lambda_{1},\dots,\lambda_{s})$ onto the $s$-dimensional downward closed simplex $\mathcal{S}_{R}:=\{{\mathbf{y}}\in\mathbb{R}^{s}~|~{\mathbf{y}}\geq 0,~\sum_{i=1}^{s}{\mathbf{y}}_{i}\leq R\}$ to obtain the new vector of eigenvalues $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$. Then, we set ${\mathbf{v}}_{t}=\sum_{i=1}^{s}\lambda_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ (i.e., the eigenvectors remain unchanged). For the correctness of this procedure see Lemma 3 in the appendix.
[…]
@@
[…]
@@
As in Remark 1, here also we can extend Problem (8) to include slightly more involved constraints of the form $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R\}$, where $\mathcal{C}\subset\mathbb{S}^{n}$ is closed and convex, as long as $\textrm{Tr}({\mathbf{X}}^{*})=R$. In particular, when $\mathcal{C}$ is a ball w.r.t. a Schatten $p$-norm ($\ell_{p}$ norm w.r.t. the eigenvalues), one can extend Lemma 3 accordingly.
4 Some Numerical Evidence
We present simple numerical simulations to demonstrate the benefit of the improved convergence rate of our method, which depends on a mixed-norm condition number, in the context of the $\ell_{1}$-constrained Problem (1). Similarly to the example in Eq. (3), we set $f({\mathbf{x}})=\frac{1}{2}({\mathbf{x}}-{\mathbf{x}}^{*})^{\top}({\mathbf{I}}+3\cdot\mathbf{1}\mathbf{1}^{\top})({\mathbf{x}}-{\mathbf{x}}^{*})$, where $\mathbf{1}$ denotes the all-ones vector. For this problem we have that the $\ell_{2}$-quadratic growth (in fact, strong convexity) constant is $\alpha_{2}=1$, and the $\ell_{2}$ and $\ell_{1}$ Lipchitz continuity constants of the gradient are $\beta_{2}=1+3n$ and $\beta_{1}=4$, respectively. We compare the convergence rate of our Algorithm 1 to a state-of-the-art Euclidean accelerated gradient method for smooth and strongly convex optimization known as V-FISTA [2], which has worst-case convergence rate $O((1-\sqrt{\alpha_{2}/\beta_{2}})^{t})$ (see Theorem 10.42 in [2]).

### v3

@@ 2 Unified Setup, Algorithm, and Convergence Analysis
Consider the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ of Algorithm 1 with a fixed step-size $\eta=\frac{\alpha_{2}}{4\beta(8s+s^{*}_{\mathcal{K}})}\in[0,1]$ and when $\gamma_{t}$ is set either via line-search or to the pre-defined value $\gamma_{t}=\eta$ for all $t$. Then,
$\forall t\geq 1:\quad f({\mathbf{x}}_{t+1})-f^{*}\leq\left({f({\mathbf{x}}_{t})-f^{*}}\right)\left({1-\frac{\alpha_{2}}{8\beta(8s+s^{*}_{\mathcal{K}})}}\right).$
Before we prove the theorem let us make an important remark.
Remark 1.
Note that with the value of $\eta$ listed in the theorem, the update for ${\mathbf{v}}_{t}$ depends on the coefficient $s\beta\eta=\frac{s\alpha_{2}}{4(8s+s_{\mathcal{K}}^{*})}$. First, this is independent of $\beta$. Second, as we shall detail in Section 3, in both the context of Problem (1) and Problem (2), we shall have $s_{\mathcal{K}}^{*}=\Theta(s)$. Thus, aside from the sparsity level $s$, this coefficient depends on a single parameter which is the quadratic growth constant $\alpha_{2}$. While setting this constant can be difficult in general, since Algorithm 1 is a descent method (as the proof will establish) there is a simple practical approach to circumvent this difficulty: on each iteration, try a few possible values, compute ${\mathbf{v}}_{t}$ for each possible value and take the one which reduces the function value the most (in particular such computations can run in parallel).
Proof of Theorem 1.
Fix some iteration $t$ of the algorithm. Whether $\gamma_{t}$ is chosen via line-search or using the pre-defined value $\eta$ stated in the theorem, the Lipschitz continuity of the gradient implies the following well known inequality:
@@
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2},$
where the second inequality is due to the triangle inequality and the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$.
@@
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+2\beta\eta^{2}s\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2},$
In the following let ${\mathbf{x}}_{t}^{*}=\arg\min_{{\mathbf{x}}\in\mathcal{X}^{*}}\|{{\mathbf{x}}-{\mathbf{x}}_{t}}\|_{2}$. Using the definition of ${\mathbf{v}}_{t}$ in the algorithm we have that
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+2\beta\eta^{2}s\|{{\mathbf{x}}_{t}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Using the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm with the assumption that ${\mathbf{x}}_{t}^{*}\in\mathcal{K}_{s}\subseteq\mathbb{E}_{s}$, and then Item 2 in Assumption 1 we have that,
$\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}\leq\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|^{2}\leq s_{\mathcal{K}}^{*}\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}.$
@@
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+2\beta\eta^{2}s\|{{\mathbf{x}}_{t}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}s^{*}_{\mathcal{K}}\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}+4\beta\eta^{2}s\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|_{2}^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(8s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2},$
where the last inequality follows from the definition of $\widehat{{\mathbf{x}}}_{t}$.
@@
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})-\eta\left({f({\mathbf{x}}_{t})-f^{*}}\right)+\frac{2\beta\eta^{2}(8s+s^{*}_{\mathcal{K}})}{\alpha_{2}}\left({f({\mathbf{x}}_{t})-f^{*}}\right).$
Subtracting $f^{*}$ from both sides and plugging-in our choice of step-size $\eta$, completes the proof.
@@ 3.1 $\ell_{1}$-ball setup
Here we instantiate our Algorithm 1 by setting $\|{\cdot}\|$ to be the $\ell_{1}$ norm $\|{\cdot}\|_{1}$ in $\mathbb{R}^{n}$ (and accordingly $\|{\cdot}\|_{*}=\|{\cdot}\|_{\infty}$). As already discussed in the Introduction, we assume there is a positive integer $s<<n$ such that every optimal solution $\mathbf{x}$$^{*}$ satisfies $\mathrm{nnz}({\mathbf{x}}^{*})\leq s$ and also $\|{{\mathbf{x}}^{*}}\|_{1}=R$ (i.e., all optimal solutions lie on a low-dimensional face of the $\ell_{1}$ ball).
We shall require the following simple lemma (a proof is given in the appendix for completeness).
@@
$\|{{\mathbf{x}}-{\mathbf{y}}}\|_{1}\leq 2\sqrt{\mathrm{nnz}({\mathbf{x}})}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
Let $s$ be such that $\mathrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\mathrm{nnz}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\|{{\mathbf{x}}}\|_{1}\leq R,~\mathrm{nnz}({\mathbf{x}})\leq s\}$, i.e., $\mathcal{K}_{s}$ is the restriction of the $\ell_{1}$ ball to $s$-sparse vectors. In light of Lemma 1 and the assumption that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=4s$.
@@
Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, by computing the Euclidean projection of the point ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta}\nabla{}f({\mathbf{x}}_{t})$ onto $\mathcal{K}_{s}$, is done by computing the Euclidean projection of the vector in $\mathbb{R}^{s}$ which corresponds to the $s$-largest (in absolute value) entries in ${\mathbf{z}}_{t}$ onto the $s$-dimensional $\ell_{1}$-ball of radius $R$, and then appending zeros (in the entries which correspond to the nsn-s lower entries in ${\mathbf{z}}_{t}$) to obtain again a point in the $n$-dimensional $\ell_{1}$-ball of radius $R$. See [3] for a proof of the correctness of this procedure.
@@
When applied to Problem (7), and assuming any optimal solution $\mathbf{x}$$^{*}$ satisfies $\mathrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside from gradient computations, the runtime of each iteration is $O(n\log{s})$. In particular, if $f$ is (convex) quadratic of the form $f({\mathbf{x}})=\frac{1}{2}{\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, the overall runtime of each iteration is $O(sn)$.
@@
It is straightforward to extend the result above to the case in which $\mathbf{x}$} is constrained to be in $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{x}}~|~\|{{\mathbf{x}}}\|_{1}\leq R\}$, where $\mathcal{C}\subset\mathbb{R}^{n}$ is closed and convex, as long as every optimal solution $\mathbf{x}$$^{*}$ satisfies $\|{{\mathbf{x}}^{*}}\|_{1}=R$. For instance when $\mathcal{C}=\{{\mathbf{x}}~|~{\mathbf{x}}\geq 0,\sum_{i=1}^{n}{\mathbf{x}}_{i}=R\}$, we have that $\mathcal{K}$ is simply a scaled simplex for which the assumption $\|{{\mathbf{x}}^{*}}\|_{1}=R$ always holds. Moreover, when $\mathcal{C}$ is closed under permutation of coordinates (e.g., an $\ell_{p}$ ball or its non-negative restriction), the sparse proximal step (computation of ${\mathbf{v}}_{t}$ in Algorithm 1) admits efficient implementation as detailed in [3].
3.2 Nuclear norm ball setup
@@
$\min_{{\mathbf{X}}\in\mathbb{R}^{m\times n}:~\|{{\mathbf{X}}}\|_{1}\leq R}f({\mathbf{X}}).$ (8)
Here we instantiate our Algorithm 1 by setting $\|{\cdot}\|$ to be the matrix nuclear norm $\|{{\mathbf{X}}}\|_{1}=\sum_{i=1}^{\min\{m,n\}}|{\sigma_{i}({\mathbf{X}})}|$, and accordingly $\|{{\mathbf{X}}}\|_{*}=\|{{\mathbf{X}}}\|_{\mathrm{op}}=\max_{i\in[\min\{m,n\}]}\sigma_{i}({\mathbf{X}})$ which is the operator norm (aka the spectral norm). Here also we let $\|{\cdot}\|_{2}$ denote the Euclidean (Frobenius) matrix norm. As already discussed in the Introduction, we assume there is a positive integer $s<<\min\{m,n\}$ such that every optimal solution $\mathbf{X}$$^{*}$ satisfies $\mathrm{rank}({\mathbf{X}}^{*})\leq s$ and also $\|{{\mathbf{X}}^{*}}\|_{1}=R$ (i.e., all optimal solutions lie on a low-dimensional face of the nuclear norm ball). The following lemma is analogous to Lemma 1. A proof is given in the appendix for completeness.
@@
Let ${\mathbf{X}},{\mathbf{Y}}\in\mathbb{R}^{m\times n}$ such that ${\mathbf{X}}\neq\mathbf{0}$, $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$. Then
$\|{\mathbf{X}}-{\mathbf{Y}}\|_{1}\leq 2\sqrt{2\mathrm{rank}({\mathbf{X}})}\,\|{\mathbf{X}}-{\mathbf{Y}}\|_{2}.$
Let $s$ be such that $\mathrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\mathrm{rank}({\mathbf{X}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\|{{\mathbf{X}}}\|_{1}\leq R,~\mathrm{rank}({\mathbf{X}})\leq s\}$, i.e., $\mathcal{K}_{s}$ is the restriction of nuclear norm ball to matrices with rank at most $s$. In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=8s$.
Here the hard-thresholding step in Algorithm 1 (computing the point $\widehat{{\mathbf{x}}}_{t}$) is done by computing the rank-$s$ approximation of the current iterate ${\mathbf{x}}_{t}$, which requires computing only the top $s$ components in the SVD of ${\mathbf{x}}_{t}$ (and discarding the rest).
Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, is done as follows: compute the $s$ top components (corresponding to the largest singular values) in the SVD of the matrix ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta}\nabla{}f({\mathbf{x}}_{t})$, and denote it as $\sum_{i=1}^{s}\sigma_{i}{\mathbf{u}}^{(i)}{\mathbf{v}}^{(i)\top}$, ${\mathbf{u}}^{(i)}\in\mathbb{R}^{m},{\mathbf{v}}^{(i)}\in\mathbb{R}^{n},i=1,\dots,s$. Then, project the vector of singular values $(\sigma_{1},\dots,\sigma_{s})$ onto the $s$-dimensional downward closed simplex $\mathcal{S}_{R}:=\{{\mathbf{y}}\in\mathbb{R}^{s}~|~{\mathbf{y}}\geq 0,~\sum_{i=1}^{s}{\mathbf{y}}_{i}\leq R\}$ to obtain the new vector of singular values $(\sigma_{1}^{\prime},\dots,\sigma_{s}^{\prime})$. Then, we set ${\mathbf{v}}_{t}=\sum_{i=1}^{s}\sigma_{i}^{\prime}{\mathbf{u}}^{(i)}{\mathbf{v}}^{(i)\top}$ (i.e., the singular vectors remain unchanged). For the correctness of this procedure see Lemma 4 in the appendix.
Overall, aside from the gradient computation, the runtime of each iteration of Algorithm 1 is dominated by computing two rank-$s$ SVDs of a $m\times n$ matrix, which is far more efficient than a full-rank SVD when $s<<\min\{m,n\}$.
@@
When applied to Problem (8), and assuming any optimal solution $\mathbf{X}$$^{*}$ satisfies $\mathrm{rank}({\mathbf{X}}^{*})\leq s<<\min\{m,n\}$ and $\|{{\mathbf{X}}^{*}}\|_{1}=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside from the gradient computations, the runtime of each iteration is dominated by two rank-$s$ SVDs of $m\times n$ matrices.
@@
As in Remark 2, here also we can extend Problem (8) to include slightly more involved constraints of the form $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\|{{\mathbf{X}}}\|_{1}\leq R\}$, where $\mathcal{C}\subset\mathbb{R}^{m\times n}$ is closed and convex, as long as $\|{{\mathbf{X}}^{*}}\|_{1}=R$ for any optimal solution $\mathbf{X}$$^{*}$. For instance, setting $m=n$ and $\mathcal{C}=\mathbb{S}^{n}_{+}$ — the positive semidefinite (PSD) cone, we can obtain the analogue of Problem (8) for optimization over the trace-bounded PSD cone. In particular, one can extend Lemma 4 (efficient sparse prox computations) accordingly.
4 Numerical Evidence